A frisbee is thrown up in the air and gains 12j of GPE.If the frisbee has a mass of 240g, calculate the maximum height of the Frisbee


1
Expert's answer
2020-09-21T08:29:15-0400

The gravitational potential energy of a frisbee

"{\rm GPE}=mgh"

Hence, the maximum height of a frisbee

"h=\frac{\rm GPE}{mg}=\rm \frac{12\: J}{9.8\: N/kg\times 0.24\: kg}=5.1\: m"